OpenStudy (anonymous):

Solve the absolute value equation |2n + 15| = 1. One solution is n = -7. What is the other solution?

OpenStudy (diyadiya):

2n+15=1 2n=1-15 n=-14/2 next one: 2n+15 = -1 2n=-1-15 2n=-16 n= -16/2 n= ?